Optical Pumping Settings

Power settings for Optical Pumping:
Fibre output measured with DET 110, 50KOhm input resistance.
Measure: 13 February 2009
repump = 1.09 V "platform"
OP/trap = 104 mV 0.6 ms step

Back to Laser Cooling and Trapping

Unless otherwise stated, the content of this page is licensed under Creative Commons Attribution-ShareAlike 3.0 License